PLEASE HELP AND SHOW WORK.... AMD IF THERE IS A CALCULATOR WAY PLEASE SHOW IT ON TI84.
If n= 150 and p (p-prime) = 0.92, construct a 95% confidence interval. Give your answers to three decimals. Sp< The video uses p (p-hat) in place of p.
Answer #1

a 95% confidence interval for p

p (1-) = (p-20.025 * ,P+20.025

= (0.92-aos * V/0.92 * 150 0.92), o.92#40ms * V/092 * 1500.92)) 0.92(1 -0.92) = (0.92-20.025 *

(0.92-1.9599би / 0.92 * ( 1-0.92) 0.92(1 -0.92) 150 0.92+1.95996* 150

(0.876585, 0.963415)